# Bidirectional Fourier-Enhanced Deep Operator Network for Spatio-Temporal Propagation in Multi-Mode Fibers

### v1

@@ Abstract
Keywords Deep Operator Networks $\cdot$ Fourier Neural Operators $\cdot$ Ultrafast photonics $\cdot$ Nonlinear fiber optics $\cdot$ Multimode fiber propagation
1 Introduction
The propagation of intense ultrashort pulses in Graded-Index (GRIN) Multi-Mode Fibers (MMFs) gives rise to some of the most complex spatio-temporal nonlinear phenomena in optics. Kerr-driven self-focusing, geometric parametric instability, massive intermodal coupling, multimode dispersion, and spatio-temporal instabilities combine to produce rapidly evolving transverse speckle patterns and temporal pulse reshaping over millimeter scale distances agrawal2013nonlinear; AGRAWAL2019309; Wright2015; krupa2016observation; 8094262; krupa2017spatiotemporal; 10.1063/1.5119434; Hansson:20. These dynamics underpin MMF-based high-power beam delivery, endoscopic imaging, spatio-temporal mode-locked lasers, and emerging neuromorphic optical computing platforms Krupa_2017; Ai:17; Morales-Delgado:15; Gusachenko2017; doi:10.1126/science.aao0831; PhysRevLett.126.093901; Cao2023; Paudel_2020; Rahmani:21. Yet full (3+1)D (3 spatial coordinates and 1 time coordinate) modeling using methods like Split-Step Fourier Method (SSFM) or Finite Difference-Beam Propagation Method (FD-BPM) is computationally expensive: needing GPU or cluster acceleration, and offer no native capability for inverse design kolesik2017computational; Brehler_2020.
In recent years, Deep Learning (DL) has emerged as a powerful alternative for accelerating the solution of complex nonlinear optical systems 8437734; Genty2021; 9596142; Salmela2021; teugin2021reusability; Chen2023; Fang:23; Hadad_2023; 10.1063/5.0169810; belonovskii2024vcsel; 10.1002/adpr.202500149; Freire:23. Physics-Informed Neural Networks (PINNs), Fourier Neural Operators (FNOs), and Deep Operator Networks (DeepONets) have all demonstrated impressive performance in simulating physical systems RAISSI2019686; Lu2021; li2021fourierneuraloperatorparametric; 10.5555/3495724.3496356. However, operator-learning frameworks, like FNOs, and DeepONets, remain largely unidirectional, and have only very recently begun to be applied to nonlinear optical systems MARGENBERG2024112725; long2025invertiblefourierneuraloperators; 9815178; 10195160.
Practical applications in multimode systems increasingly demand not only rapid forward prediction but also robust inverse retrieval of launch conditions from downstream measurements. Conventional numerical schemes lack a well-defined inverse operator and rely on slow iterative optimization or heuristic search, which become unstable in the strongly nonlinear regime 10.1364/AOP.484298; 10.1117/1.APN.2.6.066005. Operator-learning frameworks, by mapping directly between function spaces rather than finite parameter vectors, provide a natural paradigm for representing field evolution governed by nonlinear physics. FNOs along with DeepONets, are well suited to model systems mainly shaped by dispersive and diffractive effects.
@@ 1 Introduction
The following work is further organized into six sections. Section 2: Nonlinear Propagation Simulation and Data Generation outlines the GRIN MMF dataset generation pipeline, including GNLSE-based simulations, bidirectional pairing, and preparing it for training. Section 3: Modeling Nonlinear Propagation using Operator Learning Networks outlines the design, architecture, and training strategy of the bidirectional Fourier-enhanced DeepONet. Section 4: Results presents the results demonstrating the model’s accuracy in forward and inverse multimode propagation. Section 5: Discussion analyzes the physical meaning of the learned features, their relevance to fiber design, and possible avenues for improvement. Section 6: Conclusion summarizes the main findings and highlights the potential of the proposed operator-learning framework. Section 7: Experimental Methods summarizes the simulation methods and neural-network building blocks used in the study.
2 Nonlinear Propagation Simulation and Data Generation
The dataset used in this work is generated using a (3+1)D SSFM solver for the Generalized NonLinear Schrödinger Equation (GNLSE) in GRIN MMFs. The numerical implementation is adapted from the open-source code accompanying the work of Teğin et al teugin2021reusability. We gratefully acknowledge their contribution and the public availability of their simulation framework, which has significantly facilitated reproducible research in spatio-temporal multimode fibre optics.
2.1 Governing Equation
@@
The transverse window spans $54.144\,\mu\mathrm{m}\times 54.144\,\mu\mathrm{m}$ on a $64\times 64$ grid, and the temporal window has width $1.8\,\mathrm{ps}$ sampled with 1024 points. These resolutions ensure that nonlinear spatial reshaping, mode beating, and temporal dynamics are accurately resolved teugin2021reusability; Mafi:10; snyder1983optical; agrawal2010applications.
@@
2.2 Bidirectional Dataset Construction
@@ 2.3 Dataset Preparation
These preparation steps enhance the efficiency and stability of training, and improve the model’s ability to generalize to unseen conditions MAHARANA202291.
3 Modeling Nonlinear Propagation using Operator Learning Networks
Figure 1: Schematic of the proposed bidirectional Fourier-enhanced DeepONet for joint forward and inverse modeling of nonlinear pulse propagation in graded-index multimode fibers. Two dedicated branch networks separately process the input transverse intensity distribution $I(x,y;0)$ (or $I(x,y;z)$) and temporal intensity trace $I(t;0)$ (or $I(t;z)$) using spectral convolutions in their respective domains. The resulting latent representations are concatenated (denoted by $\oplus$) and combined via element-wise multiplication (denoted by $\otimes$) with the output of a trunk network that encodes the physical parameters (peak power $P_{0}$, propagation distance $z$, and a binary direction flag $d\in\{0,1\}$ indicating forward or inverse operation) through Fourier feature embeddings. A shared decoder followed by two lightweight output heads generates the predicted transverse speckle pattern $I(x,y;z)$ and temporal trace $I(t;z)$ (forward mode) or recovers the initial state $I(x,y;0)$ and $I(t;0)$ (inverse mode). The depth, width, and number of retained Fourier modes in each branch were adjusted empirically to balance expressive power and computational efficiency for the present (3+1)D multimode propagation task.
3.1 Architecture
@@
where $b_{k}(u)$ are the features learned by the branch network and $t_{k}(s)$ are the outputs of the trunk network forming the basis functions Lu2021.
@@
    Fourier Branch Network: The branch network is replaced with a spectral convolution module inspired by FNOs. The discrete input function $u(x)$ is transformed into Fourier space, filtered through learnable spectral weights, and then transformed back to produce the branch outputs $b_{k}^{\text{Fourier}}(u)$. This approach allows the network to capture long-range spatial correlations efficiently li2021fourierneuraloperatorparametric.
@@
    Fourier Trunk Network: The trunk network is augmented with a learned Fourier feature embedding. The physical parameters $s$ (here representing power, distance, and direction) are projected through a bank of sinusoidal features using a learnable matrix, producing a high-dimensional periodic encoding $t_{k}^{\text{Fourier}}(s)$. This enhances the expressiveness of the evaluation basis and enables the model to better represent oscillatory dependencies and nonlinear parameter interactions 10.5555/3495724.3496356.
@@
The bidirectional Fourier-enhanced DeepONet implemented in this work combines 2D spectral convolutions for transverse speckles, 1D spectral convolutions for temporal traces, and a Fourier-embedded trunk network to process the physical parameters (peak power, propagation distance, and direction flag). The concatenated branch and trunk features are passed through a shared decoder, which splits into two separate output heads for spatial and temporal predictions. This architecture, illustrated in Figure 1, enables simultaneous learning of forward and inverse mappings within a single network.
3.2 Training and Optimization
Figure 2: Left: Training and validation loss curves of the model, showing stable convergence and a small generalization gap at the selected checkpoint. Right: Illustration adapted from Goodfellow et al. Goodfellow-et-al-2016 showing the relationship between model capacity and generalization error, indicating the ideal stopping point prior to overfitting. The behavior observed in the left panel indicates that the model is stopped around this optimal zone.
The training procedure is designed to make use of the symmetry of the dataset while maintaining stable convergence in a high-dimensional nonlinear regression problem. By constructing matched forward and inverse pairs, as discussed in Sub-Section 2.2, the network sees both directions of the propagation operator within each batch. This removes the need for two separate models and allows a single set of parameters to learn a consistent mapping for both tasks. A binary direction flag is included as an additional input so the network can distinguish between forward prediction and inverse reconstruction without adding extra architectural complexity.
@@ 3.2 Training and Optimization
Validation loss is recorded after each epoch, and the model checkpoint with the lowest value is used for final evaluation. This ensures that the selected network achieves the best generalization performance. The training curves, as seen in Figure  2, show a smooth decrease in both training and validation loss, indicating stable convergence without signs of overfitting Goodfellow-et-al-2016; bishop2023deep.
4 Results
4.1 Forward Propagation
@@
The results show that the model predicts forward propagation accurately for both spatial and temporal intensities. As reported in Table 1, the RMSE and MAE values are low for $I(x,y;z)$ and $I(t;z)$, indicating good agreement with the ground-truth data on the test set. This suggests that the network generalizes well to unseen propagation conditions.
Table 1: Forward propagation performance on the held-out test set (normalized intensity scale).
@@ 4.2 Inverse Operation
The inverse results demonstrate that the model is also able to recover the input fields with good accuracy. As summarized in Table 2, the error values remain low for both the reconstructed spatial intensity $I(x,y;0)$ and the temporal intensity $I(t;0)$ on the held-out test set. While the inverse task is inherently more challenging than forward prediction, the obtained metrics indicate reliable reconstruction performance.
Table 2: Inverse operation performance on the held-out test set (normalized intensity scale).
@@ 6 Conclusion
7 Experimental Methods
@@
This symmetric splitting ensures second-order accuracy in the step size $dz$ and allows efficient computation using fast Fourier transforms (FFTs) for the linear operator. By iterating this procedure along $z$, the evolution of the full spatio-temporal field can be obtained agrawal2013nonlinear; teugin2021reusability.
@@
$\mathbf{s}\mapsto[\,\sin(2\pi\sum_{j}B_{jk}s_{j}),\;\cos(2\pi\sum_{j}B_{jk}s_{j})\,]\quad\text{for }k=1,\dots,D/2.$ (15)
This embedding enhances the expressiveness of the network, enabling it to capture nonlinear and high-frequency interactions in the input space 10.5555/3495724.3496356.
@@
The linear layer provides a fundamental building block for neural networks, enabling arbitrary affine mappings between vector spaces Goodfellow-et-al-2016; bishop2023deep.
Normalization layer improves numerical stability by rescaling activations using shared statistics. Given a set of activations $x_{i}$, the normalized output is
@@
where $\mu$ and $\sigma$ denote the mean and standard deviation computed over a chosen set of elements. The choice of normalization domain determines the variant: the statistics may be computed across a batch of samples (batch normalization), or across the features of a single sample (layer normalization) Goodfellow-et-al-2016; bishop2023deep.
@@
a widely adopted activation due to its simplicity, numerical stability, and effectiveness in deep architectures Goodfellow-et-al-2016; bishop2023deep.
Pooling is an operation that reduces the spatial resolution of feature maps by summarizing information within small local windows. It helps retain the most important structural patterns while decreasing dimensionality, making subsequent processing more efficient. In this work, max pooling was used, where each window is replaced by its maximum value Goodfellow-et-al-2016; bishop2023deep; 10.5555/248702.
Algorithm 1 AdamW Algorithm kingma2017adammethodstochasticoptimization; loshchilov2019decoupled
1:Given: $\alpha=10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-5}$, $\lambda\in\mathbb{R}$
2:Initialize time step $t\leftarrow 0$, parameter vector $\theta_{t=0}\in\mathbb{R}^{n}$
3:Initialize first moment vector $m_{t=0}\leftarrow 0$, second moment vector $v_{t=0}\leftarrow 0$
4:Initialize schedule multiplier $\eta_{t=0}\in\mathbb{R}$
5:repeat
6:  $t\leftarrow t+1$
7:  $\nabla f_{t}(\theta_{t-1})\leftarrow\text{SelectBatch}(\theta_{t-1})$ $\triangleright$ select batch and return the corresponding gradient
8:  $g_{t}\leftarrow\nabla f_{t}(\theta_{t-1})+\lambda\theta_{t-1}$
9:  $m_{t}\leftarrow\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$ $\triangleright$ here and below all operations are element-wise
10:  $v_{t}\leftarrow\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$
11:  $\hat{m}_{t}\leftarrow m_{t}/(1-\beta_{1}^{t})$ $\triangleright$ $\beta_{1}$ is taken to the power of $t$
12:  $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{2}^{t})$ $\triangleright$ $\beta_{2}$ is taken to the power of $t$
13:  $\eta_{t}\leftarrow\text{SetScheduleMultiplier}(t)$ $\triangleright$ can be fixed, decay, or also be used for warm restarts
14:  $\theta_{t}\leftarrow\theta_{t-1}-\eta_{t}\left(\alpha\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon)+\lambda\theta_{t-1}\right)$
15:until stopping criterion is met
16:return optimized parameters $\theta_{t}$

### v2

@@ Abstract
Index Terms:
Deep Operator Networks Fourier Neural Operators Ultrafast photonics Nonlinear fiber optics Multimode fiber propagation.
I Introduction
The propagation of intense ultrashort pulses in Graded-Index (GRIN) Multi-Mode Fibers (MMFs) gives rise to some of the most complex spatio-temporal nonlinear phenomena in optics. Kerr-driven self-focusing, geometric parametric instability, massive intermodal coupling, multimode dispersion, and spatio-temporal instabilities combine to produce rapidly evolving transverse speckle patterns and temporal pulse reshaping over millimeter scale distances [agrawal2013nonlinear, AGRAWAL2019309, Wright2015, krupa2016observation, 8094262, krupa2017spatiotemporal, 10.1063/1.5119434, Hansson:20]. These dynamics underpin MMF-based high-power beam delivery, endoscopic imaging, spatio-temporal mode-locked lasers, and emerging neuromorphic optical computing platforms [Krupa_2017, Ai:17, Morales-Delgado:15, Gusachenko2017, doi:10.1126/science.aao0831, PhysRevLett.126.093901, Cao2023, Paudel_2020, Rahmani:21]. Yet full (3+1)D (3 spatial coordinates and 1 time coordinate) modeling using methods like Split-Step Fourier Method (SSFM) or Finite Difference - Beam Propagation Method (FD-BPM) is computationally expensive: needing GPU or cluster acceleration, and offer no native capability for inverse design [kolesik2017computational, Brehler_2020].
In recent years, Deep Learning (DL) has emerged as a powerful alternative for accelerating the solution of complex nonlinear optical systems [8437734, Genty2021, 9596142, Salmela2021, teugin2021reusability, Chen2023, Fang:23, Hadad_2023, 10.1063/5.0169810, belonovskii2024vcsel, 10.1002/adpr.202500149, Freire:23]. Physics-Informed Neural Networks (PINNs), Fourier Neural Operators (FNOs), and Deep Operator Networks (DeepONets) have all demonstrated impressive performance in simulating physical systems [RAISSI2019686, Lu2021, li2021fourierneuraloperatorparametric, 10.5555/3495724.3496356]. However, operator-learning frameworks, like FNOs, and DeepONets, remain largely unidirectional, and have only very recently begun to be applied to nonlinear optical systems [MARGENBERG2024112725, long2025invertiblefourierneuraloperators, 9815178, 10195160].
Practical applications in multimode systems increasingly demand not only rapid forward prediction but also robust inverse retrieval of launch conditions from downstream measurements. Conventional numerical schemes lack a well-defined inverse operator and rely on slow iterative optimization or heuristic search, which become unstable in the strongly nonlinear regime [10.1364/AOP.484298, 10.1117/1.APN.2.6.066005]. Operator-learning frameworks, by mapping directly between function spaces rather than finite parameter vectors, provide a natural paradigm for representing field evolution governed by nonlinear physics. FNOs along with DeepONets, are well suited to model systems mainly shaped by dispersive and diffractive effects.
@@ I Introduction
The following work is further organized into six sections. Section II: Nonlinear Propagation Simulation and Data Generation outlines the GRIN MMF dataset generation pipeline, including GNLSE-based simulations, bidirectional pairing, and preparing it for training. Section III: Modeling Nonlinear Propagation using Operator Learning Networks outlines the design, architecture, and training strategy of the bidirectional Fourier-enhanced DeepONet. Section IV: Results presents the results demonstrating the model’s accuracy in forward and inverse multimode propagation. Section V: Discussion discusses the physical relevance of the learned operator and outlines directions for future improvement. Section VI: Conclusion summarizes the main findings and highlights the potential of the proposed operator-learning framework. Appendix A: Pseudo-code for the Architecture presents the algorithmic structure of the proposed model, while Appendix B: Experimental Methods describes the simulation methods and neural-network building blocks used in the study.
II Nonlinear Propagation Simulation and Data Generation
The dataset used in this work is generated using a (3+1)D SSFM solver for the Generalized NonLinear Schrödinger Equation (GNLSE) in GRIN MMFs. The numerical implementation is adapted from the open-source code accompanying the work of Teğin et al [teugin2021reusability]. We gratefully acknowledge their contribution and the public availability of their simulation framework, which has significantly facilitated reproducible research in spatio-temporal multimode fibre optics.
II-A Governing Equation
@@
The transverse window spans $54.144\,\mu\mathrm{m}\times 54.144\,\mu\mathrm{m}$ on a $64\times 64$ grid, and the temporal window has width $1.8\,\mathrm{ps}$ sampled with 1024 points. These resolutions ensure that nonlinear spatial reshaping, mode beating, and temporal dynamics are accurately resolved [teugin2021reusability, Mafi:10, snyder1983optical, agrawal2010applications].
@@
II-B Bidirectional Dataset Construction
@@ II-C Dataset Preparation
These preparation steps enhance the efficiency and stability of training, and improve the model’s ability to generalize to unseen conditions [MAHARANA202291].
Figure 1: Schematic of the proposed bidirectional Fourier-enhanced DeepONet for joint forward and inverse modeling of nonlinear pulse propagation in graded-index multimode fibers. Two dedicated branch networks separately process the input transverse intensity distribution $I(x,y;0)$ (or $I(x,y;z)$) and temporal intensity trace $I(t;0)$ (or $I(t;z)$) using spectral convolutions in their respective domains. The resulting latent representations are concatenated (denoted by $\oplus$) and combined via element-wise multiplication (denoted by $\otimes$) with the output of a trunk network that encodes the physical parameters (peak power $P_{0}$, propagation distance $z$, and a binary direction flag $d\in\{0,1\}$ indicating forward or inverse operation) through Fourier feature embeddings. A shared decoder followed by two output projections generates the predicted transverse speckle pattern $I(x,y;z)$ and temporal trace $I(t;z)$ (forward mode) or recovers the initial state $I(x,y;0)$ and $I(t;0)$ (inverse mode). The depth, width, and number of retained Fourier modes in each branch were adjusted empirically to balance expressive power and computational efficiency for the present (3+1)D multimode propagation task.
III Modeling Nonlinear Propagation using Operator Learning Networks
Figure 2: Left: Training and validation loss curves of the model, showing stable convergence and a small generalization gap at the selected checkpoint. Right: Illustration adapted from Goodfellow et al. [Goodfellow-et-al-2016] showing the relationship between model capacity and generalization error, indicating the ideal stopping point prior to overfitting. The behavior observed in the left panel indicates that the model is stopped around this optimal zone.
III-A Architecture
@@
where $b_{k}(u)$ are the features learned by the branch network and $t_{k}(s)$ are the outputs of the trunk network forming the basis functions [Lu2021].
@@
    Fourier Branch Network: The branch network is replaced with a spectral convolution module inspired by FNOs. The discrete input function $u(x)$ is transformed into Fourier space, filtered through learnable spectral weights, and then transformed back to produce the branch outputs $b_{k}^{\text{Fourier}}(u)$. This approach allows the network to capture long-range spatial correlations efficiently [li2021fourierneuraloperatorparametric].
@@
    Fourier Trunk Network: The trunk network is augmented with a learned Fourier feature embedding. The physical parameters $s$ (here representing power, distance, and direction) are projected through a bank of sinusoidal features using a learnable matrix, producing a high-dimensional periodic encoding $t_{k}^{\text{Fourier}}(s)$. This enhances the expressiveness of the evaluation basis and enables the model to better represent oscillatory dependencies and nonlinear parameter interactions [10.5555/3495724.3496356].
@@
The bidirectional Fourier-enhanced DeepONet implemented in this work combines 2D spectral convolutions for transverse speckles, 1D spectral convolutions for temporal traces, and a Fourier-embedded trunk network to process the physical parameters (peak power, propagation distance, and direction flag). The concatenated branch and trunk features are passed through a shared decoder, which splits into two separate output heads for spatial and temporal predictions. Figure 1 provides a high-level schematic of the model, while a more detailed algorithmic description is given in Appendix A, enabling simultaneous learning of forward and inverse mappings within a single network.
III-B Training and Optimization
The training procedure is designed to make use of the symmetry of the dataset while maintaining stable convergence in a high-dimensional nonlinear regression problem. By constructing matched forward and inverse pairs, as discussed in Sub-Section II-B, the network sees both directions of the propagation operator within each batch. This removes the need for two separate models and allows a single set of parameters to learn a consistent mapping for both tasks. A binary direction flag is included as an additional input so the network can distinguish between forward prediction and inverse reconstruction without adding extra architectural complexity.
@@ III-B Training and Optimization
Validation loss is recorded after each epoch, and the model checkpoint with the lowest value is used for final evaluation. This ensures that the selected network achieves the best generalization performance. The training curves, as seen in Figure  2, show a smooth decrease in both training and validation loss, indicating stable convergence without signs of overfitting [Goodfellow-et-al-2016, bishop2023deep].
IV Results
@@
IV-A Forward Propagation
The results show that the model predicts forward propagation accurately for both spatial and temporal intensities. As reported in Table I, the RMSE and MAE values are low for $I(x,y;z)$ and $I(t;z)$, indicating good agreement with the ground-truth data on the test set. This suggests that the network generalizes well to unseen propagation conditions.
TABLE I: Forward Propagation Performance on the Held-Out Test Set (Normalized Intensity Scale).
@@ IV-A Forward Propagation
IV-B Inverse Operation
The inverse results demonstrate that the model is also able to recover the input fields with good accuracy. As summarized in Table II, the error values remain low for both the reconstructed spatial intensity $I(x,y;0)$ and the temporal intensity $I(t;0)$ on the held-out test set. While the inverse task is inherently more challenging than forward prediction, the obtained metrics indicate reliable reconstruction performance.
TABLE II: Inverse Operation Performance on the Held-Out Test Set (Normalized Intensity Scale).
@@ VI Conclusion
Appendix A Pseudo-code for the Architecture
Algorithm 1 Bidirectional Fourier-Enhanced DeepONet
1:Input: Spatial field $\mathbf{u}_{s}\in\mathbb{R}^{H\times W}$, Temporal field $\mathbf{u}_{t}\in\mathbb{R}^{T}$, Physical parameters $\boldsymbol{\theta}=[P_{0},z,d]$ where $d=0$ for forward and $d=1$ for inverse operation
2:Hyperparameters: Latent dimension $p$, Fourier modes $(m_{2d},m_{1d})$, Hidden dimension $d_{h}$, Number of spectral blocks $N_{s},N_{t}$
3:1. Spatial Branch Network $\mathcal{B}_{s}$:
4:   Initialize $\mathbf{z}_{s}=\mathbf{u}_{s}$
5:   For $i=1$ to $N_{s}$:
6:    Apply SpectralConv2D($m_{2d}$) $\rightarrow$ LayerNorm $\rightarrow$ Activation $\rightarrow$ MaxPool2D
7:   Flatten and project through MLP to latent vector $\mathbf{b}_{s}\in\mathbb{R}^{p}$
8:2. Temporal Branch Network $\mathcal{B}_{t}$:
9:   Initialize $\mathbf{z}_{t}=\mathbf{u}_{t}$
10:   For $i=1$ to $N_{t}$:
11:    Apply SpectralConv1D($m_{1d}$) $\rightarrow$ LayerNorm $\rightarrow$ Activation $\rightarrow$ MaxPool1D
12:   Flatten and project through MLP to latent vector $\mathbf{b}_{t}\in\mathbb{R}^{p}$
13:3. Trunk Network $\mathcal{T}$:
14:   Apply Fourier feature embedding to $\boldsymbol{\theta}$, then MLP to $\mathbf{t}\in\mathbb{R}^{2p}$
15:4. Combine Branches and Trunk:
16:   Concatenate branch outputs: $\mathbf{b}=[\mathbf{b}_{s};\mathbf{b}_{t}]\in\mathbb{R}^{2p}$
17:   Element-wise multiply with trunk output: $\mathbf{c}=\mathbf{b}\odot\mathbf{t}$ $\triangleright$ DeepONet operator
18:5. Shared Decoder: Pass $\mathbf{c}$ through MLP decoder to hidden representation $\mathbf{h}\in\mathbb{R}^{d_{h}}$
19:6. Output Heads:
20:   Spatial output: $\hat{\mathbf{u}}_{s}=$ linear head mapping $\mathbf{h}\rightarrow\mathbb{R}^{H\times W}$
21:   Temporal output: $\hat{\mathbf{u}}_{t}=$ linear head mapping $\mathbf{h}\rightarrow\mathbb{R}^{T}$
22:return $\hat{\mathbf{u}}_{s},\hat{\mathbf{u}}_{t}$
Appendix B Experimental Methods
B-A SSFM
@@
This symmetric splitting ensures second-order accuracy in the step size $dz$ and allows efficient computation using fast Fourier transforms (FFTs) for the linear operator. By iterating this procedure along $z$, the evolution of the full spatio-temporal field can be obtained [agrawal2013nonlinear, teugin2021reusability].
B-B Spectral convolution 2D
@@ B-D Fourier Embedding
$\begin{split}\mathbf{s}\mapsto&[\,\sin(2\pi\sum_{j}B_{jk}s_{j}),\;\cos(2\pi\sum_{j}B_{jk}s_{j})\,]\\ &\quad\text{for }k=1,\dots,D/2.\end{split}$ (15)
This embedding enhances the expressiveness of the network, enabling it to capture nonlinear and high-frequency interactions in the input space [10.5555/3495724.3496356].
B-E Linear layer
@@
The linear layer provides a fundamental building block for neural networks, enabling arbitrary affine mappings between vector spaces [Goodfellow-et-al-2016, bishop2023deep].
B-F Normalization layer
Normalization layer mproves numerical stability by rescaling activations using shared statistics. Given a set of activations $x_{i}$, the normalized output is
@@ B-F Normalization layer
where $\mu$ and $\sigma$ denote the mean and standard deviation computed over a chosen set of elements. The choice of normalization domain determines the variant: the statistics may be computed across a batch of samples (batch normalization), or across the features of a single sample (layer normalization) [Goodfellow-et-al-2016, bishop2023deep].
B-G Activation function
@@
a widely adopted activation due to its simplicity, numerical stability, and effectiveness in deep architectures [Goodfellow-et-al-2016, bishop2023deep].
B-H Pooling
Pooling is an operation that reduces the spatial resolution of feature maps by summarizing information within small local windows. It helps retain the most important structural patterns while decreasing dimensionality, making subsequent processing more efficient. In this work, max pooling was used, where each window is replaced by its maximum value [Goodfellow-et-al-2016, bishop2023deep, 10.5555/248702].
B-I AdamW
Algorithm 2 AdamW Algorithm [kingma2017adammethodstochasticoptimization, loshchilov2019decoupled]
1:given $\alpha=5\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, $\lambda\in\mathbb{R}$
2:initialize time step $t\leftarrow 0$, parameter vector $\theta_{t=0}\in\mathbb{R}^{n}$, first moment vector $m_{t=0}\leftarrow\theta$, second moment vector $v_{t=0}\leftarrow\theta$, schedule multiplier $\eta_{t=0}\in\mathbb{R}$
3:repeat
4:  $t\leftarrow t+1$
5:  $\nabla f_{t}(\theta_{t-1})\leftarrow\text{SelectBatch}(\theta_{t-1})$ $\triangleright$ select batch and return the corresponding gradient
6:  $g_{t}\leftarrow\nabla f_{t}(\theta_{t-1})$
7:  $m_{t}\leftarrow\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$ $\triangleright$ here and below all operations are element-wise
8:  $v_{t}\leftarrow\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$
9:  $\hat{m}_{t}\leftarrow m_{t}/(1-\beta_{1}^{t})$ $\triangleright$ $\beta_{1}$ is taken to the power of $t$
10:  $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{2}^{t})$ $\triangleright$ $\beta_{2}$ is taken to the power of $t$
11:  $\eta_{t}\leftarrow\text{SetScheduleMultiplier}(t)$ $\triangleright$ can be fixed, decay, or also be used for warm restarts
12:  $\theta_{t}\leftarrow\theta_{t-1}-\eta_{t}\left(\alpha\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon)+\lambda\theta_{t-1}\right)$
13:until stopping criterion is met
14:return optimized parameters $\theta_{t}$
B-J Gradient clipping
Gradient clipping is a technique used during training to bound the size of gradients and prevent unstable or exploding updates. When using an $L_{2}$ norm threshold, the gradient vector $\mathbf{g}$ is rescaled if its norm exceeds a maximum value (here, 1.0). Formally, if $\|\mathbf{g}\|_{2}>1.0$, then
$\mathbf{g}\leftarrow\frac{\mathbf{g}}{\|\mathbf{g}\|_{2}}\times 1.0,$ (21)
ensuring that the $L_{2}$ norm of the gradient does not exceed 1.0 while preserving its direction. This stabilizes training by preventing excessively large parameter updates, which is especially useful during backpropagation in deep learning optimization [Goodfellow-et-al-2016, bishop2023deep, zhang2020gradientclipping].
Acknowledgment
The authors acknowledge
Conflict of Interest
The authors declare no conflict of interest.